# Unifews: Unified Entry-Wise Sparsification for Efficient Graph Neural Network

### v1

@@ 2. Related Work
GNN Propagation Personalization. To make use of the iterative model propagation and perform more adaptive maneuvers, some studies design fine-grained optimizations on each diffusion step in a node-dependent fashion to resolve GNN defects such as inefficiency (Chiang et al., 2019; Chen et al., 2020) and over-smoothing (Li et al., 2018). NDLS (Zhang et al., 2021) and NIGCN (Huang et al., 2023) personalize the hop number per node in the decoupled GNN design as an approach to acknowledge the local structure and mitigate over-smoothing. SCARA (Liao et al., 2022, 2023b) replaces the hop-base message-passing with a graph-oriented pushing scheme and eager termination, while Shadow-GNN (Zeng et al., 2021) explicitly selects the appropriate node-wise aggregation neighborhood. Another set of works (Lai et al., 2020; Spinelli et al., 2021; Miao et al., 2021; Zhang et al., 2022) achieves customized diffusion by gradually learning the hop-level importance of nodes at the price of introducing more overhead, which are less relevant to the efficient GNN computation. These personalizations can be equivalently regarded as modifications on graphs with more flexible and dedicated arrangements, as depicted in Figure 1(a). Our Unifews is akin to this idea, but with a more detailed entry-wise scheme and a better suitability for both iterative and decoupled propagations.
@@
GNN Architecture Compression. The concept of GNN pruning takes the neural network architecture into account and exploits sparsification for efficiency without hindering effectiveness (Zhang et al., 2023b). Zhou et al. (2021) consider structural channel pruning by reducing the dimension of weight matrices, while CGP (Liu et al., 2023a) appends a prune-and-regrow scheme on the graph structure. A line of research refers to the lottery ticket hypothesis (Frankle and Carbin, 2019) in the context of graph learning in search of a smaller yet effective substructure in the network. Their general pipeline, as outlined in Figure 1(b), involves a model compressor and a static graph sparsifier. Among these approaches, GEBT (You et al., 2022) finds the initial subnetwork with simple magnitude-based pruning on both adjacency and weight matrices, whilst ICPG (Sui et al., 2021) learns the importance of edges as an external task. GLT (Chen et al., 2021) and DGLT (Wang et al., 2023) employ an alternative optimization strategy that progressively processes the graph and network components. Other compression techniques such as quantization (Ding et al., 2021; Bahri et al., 2021; Wang et al., 2021b) are also investigated. We mark the difference between our work and compression methods that the latter graph and network sparsification are loosely coupled, usually entailing a training procedure on the full-size graph and model to identify sparsification components, which prevents their application to large-scale data.
3. Background
In Section 3.1, we introduce the preliminary message-passing scheme of GNN architectures by distinguishing two mainstream designs known as the iterative and decoupled GNNs. Then, we formulate the overall GNN learning process utilizing a graph smoothing framework in Section 3.2. Table 1 summarizes the primary symbols and notations used in the paper.
3.1. Graph Neural Network
Consider a self-looped graph (Kipf and Welling, 2017) $\mathcal{G}=\langle\mathcal{V},\mathcal{E}\rangle$ with $n=|\mathcal{V}|$ nodes and $m=|\mathcal{E}|$ edges. The neighborhood of a node $u\in\mathcal{V}$ is $\mathcal{N}(u)=\{v|(u,v)\in\mathcal{E}\}$, and its degree $d(u)=|\mathcal{N}(u)|$. The self-looped graph adjacency matrix is $\bar{\bm{A}}\in\mathbb{R}^{n\times n}$ and the diagonal degree matrix is $\bm{D}=\operatorname{diag}(d(u_{1}),d(u_{2}),\cdots,d(u_{n}))$. Note that diagonal entries of $\bar{\bm{A}}$ are all $1$, and self-loop edges are included in $\bm{D}$. We adopt the general graph normalization with coefficient $r\in[0,1]$, where the normalized adjacency matrix is $\tilde{\bm{A}}=\bm{D}^{r-1}\bar{\bm{A}}\bm{D}^{-r}$ and the normalized Laplacian matrix is $\tilde{\bm{L}}=\bm{I}-\tilde{\bm{A}}$.
We identify two categories of GNN architectures. An iterative GNN, represented by GCN (Kipf and Welling, 2017), takes the node attribute matrix $\bm{X}\in\mathbb{R}^{n\times f}$ as input, and recurrently computes the node representations by applying diffusion $\bm{T}$. The representation matrix $\bm{H}_{(l+1)}$ of the $(l+1)$-th layer is thus updated as:
(1) $\bm{H}_{(l+1)}=\sigma\left(\bm{T}\bm{H}_{(l)}\bm{W}_{(l)}\right),\quad l=0,1,\cdots,L-1,$
where $\sigma(\cdot)$ denotes the activation function such as ReLU or softamx. Eq. 1 implies two consecutive steps of iterative GNN updates, that the graph propagation computes the layer embedding $\bm{P}_{(l)}=\bm{T}\bm{H}_{(l)}$, and the feature transformation multiplies the learnable weight $\bm{W}_{(l)}$. For GCN, the diffusion matrix is $\bm{T}=\tilde{\bm{A}}$, whereas in the case of GAT (Veličković et al., 2018; Brody et al., 2022), the diffusion matrix $\bm{T}$ is composed of the attention weights. For simplicity, we assume the width of node representations is uniformly $f$ across all layers.
The other variant of GNN architecture, namely decoupled GNN, aims to simplify Eq. 1 by separating the propagation from iterative updates (Wu et al., 2019; Gasteiger et al., 2019; Wang et al., 2021a; Liao et al., 2023a). The graph-related propagation operations can be computed in advance as the embedding matrix $\bm{P}_{(L)}$, and the transformation is as simple as an MLP. Here we formulate the decoupled GNN with the two consecutive stages as:
(2) $\small\bm{P}_{(l+1)}$ $=\bm{T}_{(l)}\cdot\bm{P}_{(l)},\quad$ $l=0,1,\cdots,L-1,$
(3) $\bm{H}_{(l+1)}$ $=\sigma\left(\bm{H}_{(l)}\bm{W}_{(l)}\right),\quad$ $l=0,1,\cdots,L-1,$
where the boundary conditions are $\bm{P}_{(0)}=\bm{X}$ and $\bm{H}_{(0)}=\bm{P}_{(L)}$. As an exemplar, in SGC (Wu et al., 2019), there is $\bm{T}=\tilde{\bm{A}}$ and $\bm{P}_{(L)}=\tilde{\bm{A}}^{L}\bm{X}$. For personalized propagation methods (Zhang et al., 2021; Liao et al., 2022), the diffusion matrix $\bm{T}_{(l)}$ varies across different layers. Note that the hop number $L$ in Eq. 2 and Eq. 3 can be different, while we use the same notation $L$ interchangeably unless specifically mentioned.
Table 1. Summary of primary symbols and notations.

### v2

@@ 4.2 Unifews for Graph Propagation
[…]
Algorithm 1 Unifews on Decoupled Propagation
[…]
We illustrate the application of Unifews on decoupled graph propagation in Algorithm 1 as highlighted components. Note that the nested loops starting from Algorithm 1 are identical to the canonical graph propagation conducted by sparse-dense matrix multiplication in common GNNs. Compared to the normal propagation Algorithm 1, it refrains the value update and prunes the corresponding edge for insignificant entries in Algorithm 1. Hence, Unifews can be implemented into the GNN computation to enhance efficiency without additional matrix-wise overhead. For multi-dimensional feature matrix, the pruning function Equation 5 is performed by replacing $|\bm{p}[v]|$ with the specific norm $\|\bm{P}[v]\|$ across feature dimensions. Note that we adopt skip connections to each layer in the algorithm by initializing $\bm{P}_{(l+1)}$ explicitly with the value in the previous hop, in order to preserve at least one meaningful value for each node embedding in case all connected edges are pruned.
Approximation Bound.
From the theoretical perspective, the error introduced by multi-hop propagation is more complicated than the single-layer case in Theorem 4.2, as it is composed of both the diffusion and embedding values in approximation. By exploiting the GNN graph smoothing process in Definition 3.1, we show that:
$\|\hat{\bm{p}}_{(l+1)}-\bm{p}_{(l+1)}\|\leq\|\hat{\bm{p}}_{(l)}-\bm{p}_{(l)}\|+O(\epsilon)\cdot\|\hat{\bm{T}}-\bm{T}\|_{2}\|\bm{p}_{(l)}\|.$
Therefore, as long as the sparsification satisfies Theorem 4.2 for each hop, the whole process of pruning and accumulating the sparsified graph across multiple hops can be regarded as the approximate smoothing in Lemma 3.3, as stated in the following proposition and detailed in Section B.4:
Proposition 4.3 (Progressive Unifews for graph propagation).
For an $L$-hop graph propagation under Algorithm 1, if the final sparsifier satisfies $\epsilon$-similarity, then the overall process is an $\epsilon$-approximation to the original graph smoothing problem.
Moreover, the feature transformation stage in decoupled GNNs is feasible to the weight pruning described in the following section, which is similar to classical irregular pruning with magnitude-based thresholds for MLP networks (Han et al., 2015; Srinivas & Babu, 2015).
[…]
Algorithm 2 Unifews on Iterative GNN
[…]
4.3 Unifews for Iterative Update
Compared to decoupled designs, graph propagation and feature transformation in iterative GNN models are tightly integrated in each layer. Traditionally, this poses a challenge to GNN sparsification methods depicted in Figure 2(b), as specialized schemes are necessary for simultaneously modifying the two components without impairing the performance. On the contrary, our Unifews approach takes advantage of this architecture for jointly sparsifying the model towards a win-win situation in graph learning.
The sparsification of iterative Unifews for the entire message-passing process is presented in Algorithm 2, where the difference from the common scheme is also highlighted. Similarly, although presented as nested loops in the algorithm, the multiplication and sparsification are implemented as matrix operations. Its graph propagation stage is identical to Algorithm 1, except that the embedding $\hat{\bm{P}}_{(l)}$ is initialized by the previous representation $\hat{\bm{H}}_{(l)}$. Meanwhile, sparsification for weights is relatively straightforward compared to graph edges, since weight matrices are structured and their approximation is well-studied as network pruning (Han et al., 2015; Srinivas & Babu, 2015; Deng et al., 2020a). Given the embedding of the current layer $\hat{\bm{P}}_{(l)}$, we similarly rewrite the iterative GNN update as:
$\small\bm{H}_{(l+1)}[:,i]=\sigma\Big{(}\sum_{j=1}^{f}\bm{W}_{(l)}[j,i]\cdot\bm{P}_{(l)}[:,j]\Big{)},$ (8)
$\small\hat{\bm{W}}[j,i]=\operatorname{thr}_{\delta^{\prime}_{w}}\big{(}\bm{W}[j,i]\big{)}\cdot\bm{W}[j,i],\;\delta^{\prime}_{w}=\delta_{w}/\|\bm{P}[:,j]\|.$ (9)
(a) GCN Edge Pruning
(b) GCN Weight Pruning
(c) GAT Edge Pruning
(d) GAT Weight Pruning
Figure 3: Accuracy of iterative models over cora, while more results are in Figure 12. Columns of “Edge” and “Weight Sparsity” present the average results of models with solely edge and weight sparsification, respectively.
Approximation Bound.
To derive the precision bound for Unifews on iterative models, we investigate the difference of layer representation:
$\small\hat{\bm{H}}_{(l+1)}-\bm{H}_{(l+1)}=(\hat{\bm{P}}_{(l)}-\bm{P}_{(l)})\bm{W}_{(l)}+\hat{\bm{P}}_{(l)}(\hat{\bm{W}}_{(l)}-\bm{W}_{(l)}).$
The first term is similar to graph propagation in Proposition 4.3, while the second term can be bounded by the weight pruning scheme regarding $\delta_{w}$. Hence, we are able to show that the representation under layer-progressive Unifews for iterative networks containing both graph and weight operations satisfies the following proposition:
Proposition 4.4 (Progressive Unifews for iterative update).
For an $L$-round iterative update under Algorithm 2, the approximation error on output $\|\hat{\bm{H}}_{(L)}-\bm{H}_{(L)}\|_{F}$ is bounded by $O(\epsilon\|\bm{H}_{(L)}\|_{F}+\delta_{w})$.
The detailed interpretation of Proposition 4.4 is discussed in Section B.5. In brief, the approximation of layer representation is jointly bounded by factors depicting graph and weight sparsification processes. Hence, the unified pruning produces an advantageous approximation of the learned representations across GNN layers, which completes our framework for characterizing the approximation bound of Unifews for general GNN schemes by examining the graph smoothing optimization throughout model learning.
Complexity Analysis.
As shown in Algorithms 1 and 2, entry-level operations can be naturally inserted into the computation without additional overhead. For graph propagation, when the graph sparsity is $\eta_{a}=q_{a}/m$, where $q_{a}$ is the number of removed edges, the computation complexity for propagation in each layer is reduced from $O(m)$ to $O\big{(}(1-\eta_{a})m\big{)}$. In particular, for iterative models, this applies to both time and memory overhead. For weight pruning regarding matrices $\bm{P},\bm{H}\in\mathbb{R}^{n\times f}$ and $\bm{W}\in\mathbb{R}^{f\times f}$, let the pruning ratio of weight matrix be $\eta_{w}$. The sparsification scheme at least reduces complexity to $O\big{(}(1-\eta_{w})nf^{2}\big{)}$. The favorable merit of joint graph and weight sparsification by Unifews is that, both the propagation result $\bm{P}$ and the weight multiplication product $\bm{H}$ enjoy sparsity from the previous input alternatively. In fact, as proven by (Zhang et al., 2023b), the scale of reduced computational operation can be advanced to $(1-\eta_{w})^{2}$ for certain distribution.
5 Experimental Evaluation
Datasets and Metrics.
In the main experiment, we adopt 6 representative datasets including 3 small-scale (Kipf & Welling, 2017) and 3 large-scale ones (Hu et al., 2020) considering the applicability of evaluated methods. We mainly focus on the transductive node classification task. The efficiency is comprehensively assessed by computation time and floating-point operations (FLOPs), and $1$\mathrm{F}\mathrm{L}\mathrm{O}\mathrm{P}\mathrm{s}$\approx 2$\mathrm{M}\mathrm{A}\mathrm{C}\mathrm{s}$$. More datasets are evaluated in Section C.1.
Backbone Models and Sparsification Methods.
We employ GCN (Kipf & Welling, 2017) and GAT (Brody et al., 2022) as iterative backbone GNNs, since they are commonly used for sparsification. Four graph and joint compression methods are considered as baselines: GLT (Chen et al., 2021), GEBT (You et al., 2022), CGP (Liu et al., 2023a), and DSpar (Liu et al., 2023b). Additionally, we examine the random entry-wise scheme RAN by randomly removing entries according to the given sparsity. Backbones of decoupled models are SGC (Wu et al., 2019) and APPNP (Klicpera et al., 2019). Two baselines with propagation personalization techniques, NDLS (Zhang et al., 2021) and NIGCN (Huang et al., 2023), are used, while both are only available for SGC.
Hyperparameters.
We commonly utilize graph normalization $r=0.5$, model layer depth $L=2$, and layer width $f_{hidden}=512$. For decoupled models, the number of propagation hops is $20$. We employ full-batch and mini-batch training for iterative and decoupled methods, respectively. The total number of training epochs is $200$, including pre-training or fine-tuning process in applicable methods. The batch size is $512$ for small datasets and $16384$ for large ones. We tune the edge and weight sparsity of evaluated models across their entire available ranges.
5.1 Performance Comparison
We first separately apply only one part of sparsification, i.e., either edge or weight pruning, for better comparison. Figure 3 presents accuracy results of iterative backbones and compression methods over representative datasets. GEBT encounters out of memory error on pubmed. For larger graphs, most of the iterative baselines suffer from the out of memory error due to the expense of trainable adjacency matrix design and full-graph training process.
Table 1: Results of $\eta_{a}=50\%$ graph sparsification on decoupled models. “Time” denotes propagation time (in seconds), while “FLOPs” separately presents the operations (in $\mathrm{G}\mathrm{M}\mathrm{A}\mathrm{C}\mathrm{s}$) for propagation and transformation on all nodes. “Improvement” is the comparison between Unifews and the backbone model, where improved accuracy is marked in bold fonts.
Dataset cora citeseer pubmed arxiv products papers100m Nodes $n$ $2,485$ $3,327$ $19,717$ $169,343$ $2,400,608$ $111,059,956$ Edges $m$ $12,623$ $9,228$ $88,648$ $2,315,598$ $123,718,024$ $3,228,124,712$ Metric Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs Acc Time FLOPs SGC 85.8 0.13 0.36+2.2 67.7 0.44 0.68+2.8 83.0 0.36 0.89+2.3 68.8 3.9 5.9+15.0 79.1 289.6 247.4+434.4 63.3 19212 17.7+253.6 +NDLS 80.3 1362 0.19+2.7 64.7 1940 0.33+4.3 77.0 4717 0.42+2.0 (OOM) (OOM) (OOM) +NIGCN 84.2 0.45 0.22+3.0 70.4 0.47 0.28+2.1 85.0 9.2 0.44+2.0 63.7 87.6 15.6+14.7 77.9 1026 137.2+182.0 53.7 1770 110.7+238.6 +Unifews 86.0 0.10 0.18+1.2 73.0 0.26 0.35+1.7 83.0 0.24 0.47+2.0 69.4 1.5 3.0+13.3 78.5 203.1 124.0+186.9 63.1 192.4 5.3+143.8 Improvement 0.2 1.3× 2.0×, 1.9× 5.3 1.7× 2.0×, 1.7× 0.0 1.5× 1.9×, 1.2× 0.6 2.6× 2.0×, 1.1× -0.5 1.4× 2.0×, 2.3× -0.2 99.8× 3.3×, 1.8× APPNP 86.2 0.15 0.36+2.2 71.6 0.43 0.68+2.8 87.6 0.33 0.89+2.3 64.8 2.6 5.9+20.9 72.5 248.5 247.4+269.4 60.9 15305 17.7+247.7 +Unifews 86.5 0.08 0.18+1.8 73.7 0.21 0.31+2.3 88.0 0.26 0.43+1.8 65.0 0.93 3.0+15.0 76.9 58.5 31.7+186.9 62.8 178.5 8.9+241.8 Improvement 0.4 1.8× 2.0×, 1.2× 2.1 2.0× 2.2×, 1.2× 0.4 1.3× 2.1×, 1.3× 0.2 2.8× 1.9×, 1.4× 4.5 4.2× 7.8×, 1.4× 1.9 85.7× 2.0×, 1.0×
Edge Sparsification.
For the iterative architecture in Figures 3(a) and 3(c), Unifews outperforms state-of-the-art graph and joint compression approaches in most backbone-dataset combinations. Typically, for relatively small ratios $\eta_{a}<80\%$, models with Unifews pruning achieve comparable or exceeding accuracy, aligning with our approximation analysis. For higher sparsity, Unifews benefits from the skip connection design, which carries essential information of node identity and therefore retains accuracy no worse than the trivial transformation. On the contrary, most of the competitors experience significant accuracy drop on one or more datasets. CGP and DSpar exhibit poor utility on the GAT backbone since its variable connections are more vulnerable to removal. Additionally, the comparison with random sparsification indicates that, the entry-wise scheme is particularly effective for small ratios, as the randomized pruning even surpasses dedicated methods for some circumstances. For high edge sparsity, the Unifews preservation of dominant edges and identity mapping is critical to accuracy, which validates the advantage of our design.
Similarly, for decoupled propagation in Figure 4, Unifews is able to preserve remarkable accuracy and outperform personalized propagation methods. On citeseer, it raises SGC accuracy by $2\%$ through mitigating the over-smoothing issue. We hence conclude that, compared to heuristic pruning schemes, Unifews successfully removes unimportant graph edges without sacrificing efficacy, thanks to its fine-grained and adaptive scheme.
(a) Edge Pruning on cora
(b) Edge Pruning on citeseer
Figure 4: Accuracy of decoupled models over cora and citeseer. Unifews is employed with solely edge removal.
Weight Sparsification.
Figures 3(b) and 3(d) display the efficacy of network pruning. For all combinations of models and graphs, Unifews achieves top-tier accuracy with a wide range of weight sparsity. For small ratios, the model produces up to $3\%$ accuracy gain over the bare GNN, resembling the benefit of neural network compression (Hoefler et al., 2021). Most evaluated baselines also maintain low error rate compared to their performance in graph pruning, indicating that the GNN weights are relatively redundant and are suitable for substantial compression. Thanks to the redundancy in GNN architecture, baselines including GEBT and CGP present strong performance in certain cases compared to edge sparsification. Random removal fails for high weight sparsity, which implies that the magnitude-based scheme is the key to maintain model effectiveness.
(a) Edge Pruning, $\eta_{w}=30\%$
(b) Weight Pruning, $\eta_{a}=30\%$
Figure 5: Accuracy of joint sparsification with varying edge and weight ratios, while fixing the other at $30\%$.
Joint Sparsification.
For unified pruning on graph edges and network weights, accuracy comparison is provided by Figure 5 with applicable methods. It is noticeable that Unifews retains comparable or better accuracy than the backbone GCN with up to $3\%$ improvement. Most baseline methods only obtain suboptimal accuracy especially for weight pruning, which is affected by their comparatively poor graph sparsification. While GEBT mostly retains effectiveness on cora, it is highly limited by the specific architecture and high training overhead. The evaluation further highlights the advantage of Unifews in considering the two GNN operation stages in a unified manner.
5.2 Efficiency Analysis
FLOPs.
As graph computation is the primary bottleneck for large-scale tasks, we particularly study the efficiency improvement utilizing decoupled models in Table 1. Evaluation implies that, Unifews is effective in producing operational reduction proportional to the sparsity, i.e., saving $50\%$ computation FLOPs. It also achieves higher compression such as for APPNP over products, by recognizing more unnecessary edges than required. Contrarily, NDLS suffers from out of memory error on large datasets due to its complex calculation and implementation, while NIGCN does not guarantee decreased computation because of its expansionary propagation design. Table 1 also presents the transformation FLOPs for reference, where the sparsified embedding of Unifews also benefits the downstream computation. We remark that, although the FLOPs value for transformation stage appears to be on the same level with propagation, in practice it can be efficiently processed utilizing parallel computation on devices such as GPU (Liu et al., 2020; Peng et al., 2022). Hence the GNN scalability bottleneck, especially on large datasets, is still the graph propagation. In this context, the ability of Unifews in reducing propagation operations is of unique importance.
(a) GCN on arxiv
(b) GAT on arxiv
(c) SGC on arxiv
(d) SGC on products
Figure 6: Propagation and transformation FLOPs. Iterative models: (a) GCN and (b) GAT over arxiv. Decoupled models: SGC over (c) arxiv and (d) products.
Runtime.
Within each method in Table 1, the propagation time is correlated with its FLOPs, and escalates rapidly with the data size due to the nature of graph operations. Comparison across methods demonstrates that Unifews excels in efficiently conducting graph propagation with regard to both FLOPs and execution time. It realizes significant acceleration on large graphs by favorably reducing the graph scale across all layers, which benefits the system workload such as better entry-level access. The papers100m result highlights the superiority of Unifews with 85100×85-100$\times$85 - 100 × improvement over the backbone and 9×9$\times$9 × speed-up over NIGCN.
Operation Breakdown.
To specifically evaluate the efficiency enhancement, in Figure 6, we separately assess the FLOPs related to graph propagation and network transformation for different backbone models and datasets under Unifews sparsification. For better presentation, model width is set to 64 in this experiment. Figures 6(a) and 6(b) imply that the majority of computational overhead of iterative models is network transformation, even on graphs as large as arxiv. Consequently, weight compression is essential for reducing GNN operations.
On the other hand, graph propagation becomes the bottleneck in decoupled designs, and is increasingly significant on graphs with greater scales. This is because of the larger number of propagation hops $L=20$ for these structures. In this case, Unifews is effective in saving computational cost by simplifying the propagation and bypassing unnecessary operations, with over 20×20$\times$20 × and 5×5$\times$5 × joint reduction on arxiv and products, respectively. We also discover the benefit of joint pruning that a higher threshold results in smaller FLOPs of both operations in Figures 6(c) and 6(d), which signifies the win-win situation brought by increased sparsity. By combining these two sparsifications, we summarize that the unified scheme of Unifews is capable of mitigating the computational overhead for both propagation- and transformation-heavy tasks.
(a) Message $\bm{\tau}_{(0)}$
(b) Representation $\bm{H}_{(1)}$
Figure 7: Entry sparsity under joint sparsification, evaluated by matrix norm relative to the unpruned one on a 2-layer GCN over cora. (a) Edge-wise message $\|\hat{\bm{\tau}}_{(0)}\|/\|\bm{\tau}_{(0)}\|$. (b) Node-wise representation $\|\hat{\bm{H}}_{(1)}\|_{F}/\|\bm{H}_{(1)}\|_{F}$.
5.3 Parameter Discussion
Sparsity.
Figure 7 displays the relative density of the edge message and node representation matrices, respectively. It can be observed that the entry sparsity is enhanced with the increase of both two pruning ratios, signifying our theoretical analysis that Unifews not only directly shrinks edges and weights, but promotes the sparsity of the product matrix as well. It also supports our claim that Unifews is superior in employing dual sparsification where the two stages benefit each other alternatively and enjoy a win-win situation brought by the increased sparsity.
(a) GCN on cora
(b) GAT on cora
(c) SGC on cora
(d) APPNP on cora
Figure 8: Sensitivity of joint sparsification thresholds on four models over cora. We respectively set the weight ratio $\delta_{w}$ to 00 and $0.5$, then evaluate the accuracy, edge sparsity, and weight sparsity of the model. Note that the x-axis representing diffusion threshold is on a logarithmic scale.
Joint Sparsification Thresholds.
The thresholds controlling diffusion and weight sparsification are pivotal to Unifews compression, affecting both efficacy and efficiency of the produced model. Figure 8 presents the changes of inference accuracy and dual sparsity of GNN models under graph and joint sparsification with varying adjacency thresholds $\delta_{a}$. Accuracy in the plot follows the conclusion in previous evaluation, that it only degrades above extreme sparsity $\eta_{a}>95\%$. The experiment reveals that, the relation between edge sparsity and adjacency threshold aligns with Theorem 4.2, and decoupled models typically require a larger range of threshold to traverse the sparsity range, which is because of the wider distribution of their entry values throughout the deeper propagation. Interestingly, the weight sparsity when $\delta_{w}=0.5$ also increases under high edge pruning ratios. The pattern is also observed in the reverse case. We deduce that the reciprocal enhancement in graph and model sparsity is brought by the unified sparsification of Unifews, which conforms to our analysis as well as previous studies (Zhang et al., 2023b).
6 Conclusion
In this work, we present Unifews, an entry-wise GNN sparsification with a unified framework for graph edges and model weights. By bridging spectral graph smoothing and GNN sparsification, we showcase in theory that the layer-progressive Unifews provides an effective approximation on the graph learning process with a close optimization objective, which is favorable for multi-layer GNN updates in both iterative and decoupled architectures. Comprehensive experiments underscore the superiority of Unifews in terms of efficacy and efficiency, including comparable or improved accuracy, $90-95\%$ operational reduction, and up to 100×100$\times$100 × faster computation.
Limitation.
One potential limitation lies in the consideration of graph heterophily, as messages with large magnitudes may not be beneficial to model prediction. In this case, the graph smoothing objective in Definition 3.1 needs to be refined; consequently, the sparsification strategy should be adjusted.
Acknowledgements
This research is supported by Singapore MOE AcRF Tier-2 Grant (T2EP20122-0003) and NTU-NAP startup grant (022029-00001). Ningyi Liao is supported by the Joint NTU-WeBank Research Centre on FinTech, Nanyang Technological University, Singapore.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Detailed Related Works and Comparison
A.1 GNNs Simplification
Iterative Graph Sparsification.
Simplification techniques mainly aim to reduce GNN message-passing operations while maintaining its scheme. Consequently, these methods usually enjoy better expressiveness and generality to different model architectures. Among them, a large portion of works sparsify edge connections for accuracy improvement, typically through deleting edges. NeuralSparse (Zheng et al., 2020) and LSP (Kosman et al., 2022) drop edges based on $k$-neighbor metric for each node and layer-wise local graph topology, respectively, while AdaptiveGCN (Li et al., 2021) and SGCN (Li et al., 2022) choose to implement sparsity by learnable edge features and graph adjacency.
In comparison, FastGAT (Srinivasa et al., 2020) and DSpar (Liu et al., 2023b) relate graph modification to spectral sparsification mainly as an attempt to boost efficiency. FastGAT eliminates connections in the GAT (Veličković et al., 2018) model by calculating the Effective Resistance (Spielman & Srivastava, 2011) but at the expense of high computational time. DSpar employs a loose approximation and successfully applies to larger datasets.
Most of the simplification techniques, however, conduct graph modification in a one-time and layer-agnostic manner, resulting in limited flexibility and utility under the trade-off between efficiency and efficacy.
Decoupled Propagation Personalization.
To make use of the model propagation and perform more adaptive maneuvers, some studies design fine-grained optimizations on each diffusion step in a node-dependent fashion to resolve GNN defects such as inefficiency (Chiang et al., 2019; Chen et al., 2020; Liu et al., 2025; Liao et al., 2026) and over-smoothing (Li et al., 2018; Liao et al., 2025). NDLS (Zhang et al., 2021) and NIGCN (Huang et al., 2023) personalize the hop number per node in the decoupled GNN design as an approach to acknowledge the local structure and mitigate over-smoothing. SCARA (Liao et al., 2022, 2023b) and LD2 (Liao et al., 2023a) replace the hop-base message-passing with a graph-oriented pushing scheme and eager termination, while Shadow-GNN (Zeng et al., 2021) explicitly selects the appropriate node-wise aggregation neighborhood.
Another set of works (Lai et al., 2020; Spinelli et al., 2021; Miao et al., 2021; Zhang et al., 2022) achieves customized diffusion by gradually learning the hop-level importance of nodes at the price of introducing more overhead, which are less relevant to the efficient GNN computation.
Table 2: Summary of primary symbols and notations.